GTS meeting 5 3 16

From Ciswikidb
Jump to: navigation, search


  • Beamline
    • Viewers are functional with EPICS control, but need analog video cameras. J. Gubeli working on this
    • Vacuum UHV signals from all six pumps are now in EPICS
    • Magnets have been wired up, control EPICS screen is done. Trim PS are being 'conditioned'. Resistor box for RadiaBeam steerers is being built. E. Diggs expects to complete all hardware work by the end of this week. Then next week DC power groups needs to perform hot checkout, polarity checks, etc before Sarin signs off on the system.
    • Dump. Bubba is piping the LCW line. Need a flowmeter from P. Francis with electrical output for MPS.
    • Checkout the screens here: [1]
  • Laser
    • Shukui has aligned the laser to the gun chamber entrance viewport!
    • The variable attenuator working now, in this case an ND with fixed OD is not needed unless you are putting it into the viewer assembly.
    • The cables for pico-motors, for laser camera, for power meter signals, seem to be missing/used by others, it'll be great if Jim's crew can help to sort out.
    • The GTS drive laser control is not working except the motor control. Shukui talked with Wes last week, he agreed to help with the GTS DL EPICS together with one of Pam's group members.
  • Cathode
    • Mamun has made a cathode, this time on the puck!
    • The QE is below 1%, but will be ideal for first beam!
  • Other..
    • There is no computer with EPICS in the GET cave, which is not really convenient when checking things out there.
  • Drawings of the new puck for magnetized beam:
media:haysg_JL0035765-LDRD_PUCK_STEEL_PLUG_ASSY.pdf
media:haysg_JL0035913-LDRD_PUCK_BODY_NEW.pdf
media:haysg_JL0035916-LDRD_PUCK_PLUG_BODY.pdf